# An Adaptive Learning Approach to Multivariate Time Forecasting in Industrial Processes

### v2

@@ 2 Time losses in industrial processes
Note that each of these categories could further be subdivided based on the specific cause of the loss, although such a classification is typically customized according to the particular nature of the process under consideration. By taking the length of an observation period as a reference -hereinafter referred to as Opening Time or OT- one can derive different production times by successively subtracting each time loss, as illustrated in Figure 2.1 and definitions (2).
[…]
Moreover, the ratio between the production times can be used to define some well-known effectiveness indicators, which are enumerated in formulae (2):
[…]
@@
[…]
@@
date date yyyy-mm-dd
start timestamp hh:mm:ss
@@ 3.2 The discrete process
$\bm{\pi}^{(\mathbf{s})}=\left[\pi_{1}^{(\mathbf{s})}\ldots\ \pi_{K}^{(\mathbf{s})}\right]$ $\sim\textrm{Dirichlet}\left(\mathbf{a}^{(\mathbf{s})}=[a^{(\mathbf{s})}_{1}\ \ldots\ a^{(\mathbf{s})}_{K}]\right)$
$\mathbf{p}_{k}^{(\mathbf{s})}=\left[p_{k1}^{(\mathbf{s})}\ \ldots\ p_{kK}^{(\mathbf{s})}\right]$ $\sim\textrm{Dirichlet}\left(\bm{\alpha}^{(\mathbf{s})}_{k}=[\alpha^{(\mathbf{s})}_{k1}\ \ldots\ \alpha^{(\mathbf{s})}_{kK}]\right).$ (3.1)
@@ 3.3 The continuous process
$\mathbf{y}_{n}|\mathbf{w}_{n}\sim N_{m}\left(\mathbf{u}_{n}\mathbf{H}_{u},\bm{\Sigma}_{u}\right)$ (3.2a)
$\mathbf{y}_{n}|c_{n}\sim N_{m}(\mathbf{v}_{n}\mathbf{H}_{v},\bm{\Sigma}_{v}),$ (3.2b)
where $m$ is the number of response variables; $\mathbf{H}_{u},\mathbf{H}_{v}$ denote coefficient matrices; $\bm{\Sigma}_{u},\bm{\Sigma}_{v}$ are covariance matrices; and $\mathbf{u}_{n}=\left[1\ \mathbf{w}_{n}\right]$, $\mathbf{v}_{n}=v(c_{n})$ are the covariate vectors at the $n$-th step, with $v(\cdot)$ a function of the hidden state. In particular, we propose considering the conditional expectation $v(c_{n})=\mathbb{E}\left[c_{n}|c_{n-1}\right]$, which, with the Dirichlet distribution assumptions, simplifies to the Dirichlet parameters in (3.1) normalized by their concentration parameters, i.e., $\mathbf{v}_{n}=\mathbf{a}/a_{0}$ or $\mathbf{v}_{n}=\bm{\alpha}_{c_{n-1}}/\alpha_{c_{n-1},0}$.
@@ 3.4 The adaptive algorithm
We explain the adaptive algorithm for the case of the parameters in distribution (3.2a); the same procedure is applied to the parameters in (3.2b). Let $\lambda\in(0,1]$ be a forgetting factor that accounts for the weight of past observations. As soon as a new sample $\mathbf{y}_{n}$ becomes available, the prior estimators $\mathbf{H}_{n-1},\bm{\Sigma}_{n-1}$ are updated to $\mathbf{H}_{n},\bm{\Sigma}_{n}$ (we omit the model subscripts in the parameters for clarity) through an adaptive algorithm described by the following equations:
$\gamma_{n}$ $=1+\lambda\gamma_{n-1}$ (3.3a)
$\mathbf{H}_{n}$ $=\mathbf{H}_{n-1}+\dfrac{\mathbf{P}_{n-1}\mathbf{u}_{n}^{T}}{\lambda+\mathbf{u}_{n}\mathbf{P}_{n-1}\mathbf{u}_{n}^{T}}\left(\mathbf{y}_{n}-\mathbf{u}_{n}\mathbf{H}_{n-1}\right)$ (3.3b)
$\bm{\Sigma}_{n}$ $=\bm{\Sigma}_{n-1}-\dfrac{1}{\gamma_{n}}\left[\bm{\Sigma}_{n-1}-\dfrac{\lambda\left(\mathbf{y}_{n}-\mathbf{u}_{n}\mathbf{H}_{n-1}\right)^{T}\left(\mathbf{y}_{n}-\mathbf{u}_{n}\mathbf{H}_{n-1}\right)}{\lambda+\mathbf{u}_{n}\mathbf{P}_{n-1}\mathbf{u}_{n}^{T}}\right]$ (3.3c)
$\mathbf{P}_{n}$ $=\dfrac{1}{\lambda}\left(\mathbf{P}_{n-1}-\dfrac{\mathbf{P}_{n-1}\mathbf{u}_{n}^{T}\mathbf{u}_{n}\mathbf{P}_{n-1}}{\lambda+\mathbf{u}_{n}\mathbf{P}_{n-1}\mathbf{u}_{n}^{T}}\right)$ (3.3d)
initialized with $\mathbf{H}_{0}=\bm{0},\ \bm{\Sigma}_{0}=\bm{0},\ \mathbf{P}_{0}=\mathbf{I}$ and $\gamma_{0}=0$.$\gamma_{n}$ is the total weight of the sample. Note that if $\lambda=1$ then all the observations have the same weight and $\gamma_{n}=n$ is the sample size; if $\lambda<1$, past data gradually loses influence over time while recent data has a greater impact. Equations (3.3a)-(3.3d) can be obtained by extending to the multivariate case the Maximum-Likelihood-based proof provided in Alvarez et al.Alvarez et al. [2021], Theorem. 1. We refer to Appendix A for an alternative proof using a Bayesian approach.
@@ 3.5 Forecasting
After the training step, each distribution (3.2a)-(3.2b) produces a forecast of the responses. Later, these forecasts are combined using a minimum-variance criterion to obtain the final prediction [Roccazzella et al., 2022]. Once a new observation is available the update-prediction loop begins again. In particular, when the parameters are updated after the $n$-th observation is received we can write
@@
$\mathbb{E}_{n}\left[\mathbf{y}_{n+1}\right]\stackrel{{\scriptstyle def}}{{=}}\hat{\mathbf{y}}_{n+1}$ $=\mathbf{u}_{n+1}\mathbf{H}_{u}\mathbf{D}+\mathbf{v}_{n+1}\mathbf{H}_{v}(\mathbf{I}-\mathbf{D})$ (3.4a)
$\mathbb{V}\mathrm{ar}_{n}\left[\mathbf{y}_{n+1}\right]\stackrel{{\scriptstyle def}}{{=}}\hat{\bm{\Sigma}}_{n+1}$ $=\mathbf{D}\bm{\Sigma}_{u}\mathbf{D}+(\mathbf{I}-\mathbf{D})\bm{\Sigma}_{v}(\mathbf{I}-\mathbf{D}),$ (3.4b)
@@
$\delta_{j}=\dfrac{\sigma_{v,j}^{2}}{\sigma_{u,j}^{2}+\sigma_{v,j}^{2}},$ (3.5)
@@ 4.3 Clustering
Once the classification is complete, each sequence in the training set is segmented into labelled intervals corresponding to the detected operational states. These states are later used to learn the parameters of the discrete part of the model, as described in Section 3.2, starting from non-informative Jeffreys’ priors for the Dirichlet distributions (3.1), that is, $a_{k}=\alpha_{jk}=1/2,\ \forall j,k$.
@@ 4.4 Pseudo-code
$\bm{\Pi}$ =$=\left\{\mathbf{a_{s}}=[a^{(\mathbf{s})}_{1}\ \ldots\ a^{(\mathbf{s})}_{K}],\mathbf{A_{s}}=\left(\alpha_{kj}^{(\mathbf{s})}\right)_{k,j\in\mathcal{K}},\ \mathbf{s}\in\mathcal{S}\right\}$
$\bm{\Psi}$ =$=\left\{\mathbf{H}_{u}^{(\mathbf{s})},\bm{\Sigma}_{u}^{(\mathbf{s})},\mathbf{H}_{v}^{(\mathbf{s})},\bm{\Sigma}_{v}^{(\mathbf{s})},\ \mathbf{s}\in\mathcal{S}\right\}$
$\bm{\Omega}$ =$=\left\{\mathbf{P}_{u}^{(\mathbf{s})},\gamma_{u}^{(\mathbf{s})},\mathbf{P}_{v}^{(\mathbf{s})},\gamma_{v}^{(\mathbf{s})},\ \mathbf{s}\in\mathcal{S}\right\}$
@@
2:   $\bm{\Pi}$triangleright Dirichlet parameters
3:   $\bm{\Psi}$triangleright model parameters
4:   $\bm{\Omega}$triangleright state parameters
5:   $\lambda_{u},\lambda_{v}$triangleright forgetting factors
6:   $\mathbf{x}_{n}=\left[\mathbf{z}_{n}\ \mathbf{w}_{n}\right]$triangleright covariates
7:   $c_{n-1},c_{n},\mathbf{y}_{n}$triangleright state labels, responses
@@
16:     Update $\gamma_{u}^{(\mathbf{s})},\mathbf{H}_{u}^{(\mathbf{s})},\bm{\Sigma}_{u}^{(\mathbf{s})},\mathbf{P}_{u}^{(\mathbf{s})}$ using equations (3.3a), (3.3b), (3.3c), (3.3d) respectively
17:     Update $\gamma_{v}^{(\mathbf{s})},\mathbf{H}_{v}^{(\mathbf{s})},\bm{\Sigma}_{v}^{(\mathbf{s})},\mathbf{P}_{v}^{(\mathbf{s})}$ using equations (3.3a), (3.3b), (3.3c), (3.3d) respectively
@@
2:   $\bm{\Psi}$triangleright model parameters
3:   $\left\{\mathbf{o}_{1},\ldots,\mathbf{o}_{K}\right\}$triangleright centroids
4:   $\mathbf{t}_{n}$triangleright classification variables
5:   $\mathbf{x}_{n+1}=\left[\mathbf{z}_{n+1}\ \mathbf{w}_{n+1}\right]$triangleright covariates
@@
7:$\hat{\mathbf{y}}_{n+1},\ \hat{\bm{\Sigma}}_{n+1}$triangleright responses forecast, prediction error
8:$k\leftarrow\operatorname*{arg\,min}_{j\in\{1,\ldots,K\}}d(\mathbf{o}_{j},\mathbf{t}_{n})$
@@
10:$\mathbf{s}\leftarrow\mathbf{z}_{n+1}$
11:$\mathbf{b}\leftarrow\begin{cases}\mathbf{a}^{(\mathbf{s})}&\text{if the sequence begins}\\ \bm{\alpha}^{(\mathbf{s})}_{k}&\text{otherwise}\end{cases}$
12:$\mathbf{u}\leftarrow\left[1\ \mathbf{w}_{n+1}\right]$
13:$\mathbf{v}\leftarrow\mathbf{b}/\left(\mathbf{b}\cdot\mathbf{1}\right)$
14:Compute $\mathbf{D}$ using equation (3.5)
15:Compute $\hat{\mathbf{y}}_{n+1}$ using equation (3.4a)
16:Compute $\hat{\bm{\Sigma}}_{n+1}$ using equation (3.4b)
@@ 5 Application to a real case study
The time variables OpT and NOpT are selected as response variables due to their strong correlation (0.83), which fully justifies the use of a multivariate model. According to the scheme presented in Figure 2.1 and the definitions (2), (2) and (2.3), predicting these variables allows for the computation of time losses due to low production speed and the performance index. Additionally, in scenarios where time measurements are equispaced and scheduled stops are predetermined (i.e., with deterministic OT, SBT and LT), these predictions also enable the estimation of time losses due to unexpected stops and the availability index.
@@ Appendix A Appendix
    $\tilde{\beta}$ $=\textrm{vec}\left(\tilde{B}\right)$
@@
    $N\tilde{S}$ $=\left(Y-X\tilde{B}\right)^{T}\left(Y-X\tilde{B}\right)+\left(\tilde{B}-B_{0}\right)^{T}V_{0}\left(\tilde{B}-B_{0}\right).$ (A.1a)
@@
    $\mathbb{E}\left[B|\Lambda\right]$ $=\tilde{B}=B_{0}+\left(X^{T}X+V_{0}\right)^{-1}X^{T}\left(Y-XB_{0}\right)$ (A.2a)
[…]
    $=\Lambda_{0}-\dfrac{N\left(\Lambda_{0}-\tilde{S}\right)}{N_{0}+N}.$ (A.2b)
@@
    $Y$ $=\ \mathbf{y}_{n}\ \text{ (new responses, row vector $1\times m$)}$
    $X$ $=\ \mathbf{u}_{n}\ \text{ (new predictors, row vector $1\times p$)}$
    $\tilde{B}$ $=\ \mathbf{H}_{n}\ \text{ (posterior mean, $p\times m$)}$
    $B_{0}$ $=\ \mathbf{H}_{n-1}\ \text{ (prior mean, $p\times m$)}$
    $V_{0}$ $=\ \lambda\mathbf{P}_{n-1}^{-1}\ \text{ (prior state matrix, $p\times p$)}$
    $(X^{T}X+V_{0})^{-1}$ $=\mathbf{P}_{n}\ \text{ (posterior state matrix, $p\times p$)}$
@@
    Equation (3.3a) is obvious given that $\gamma_{n}=1+\lambda\left(1+\lambda+\cdots\lambda^{n-2}\right)=1+\lambda\gamma_{n-1}$. From the above definition of the posterior state matrix $\mathbf{P}_{n}$ and using the Sherman-Morrison formula111$(A+uv^{T})^{-1}=A^{-1}-\dfrac{A^{-1}uv^{T}A^{-1}}{1+v^{T}A^{-1}u}$,
@@
    we obtain equation (3.3d). For the posterior mean $\tilde{B}=\mathbf{H}_{n}$, using expression (A.2a) we have
@@
    yielding equation (3.3b). On the other hand, from the definition (A.1a) we get
@@
    $N_{0}$ $=\ \lambda\gamma_{n-1}$
    $N$ = $=\ 1$
    $\Lambda_{0}$ $=\bm{\Sigma}_{n-1}\ \text{ (prior mean variance, $m\times m$)}$
    $\mathbb{E}\left[\Lambda|Y\right]$ $=\ \bm{\Sigma}_{n}\quad\text{ (posterior mean variance, $m\times m$)}$
    and using expression (A.2b)
    $\bm{\Sigma}_{n}=\mathbb{E}\left[\Lambda|Y\right]$ $=\Lambda_{0}-\dfrac{N\left(\Lambda_{0}-\tilde{S}\right)}{N_{0}+N}$

### v3

@@ 2 Time losses in industrial processes
Note that each of these categories could further be subdivided based on the specific cause of the loss, although such a classification is typically customized according to the particular nature of the process under consideration. By taking the length of an observation period as a reference -hereinafter referred to as Opening Time or OT- one can derive different production times by successively subtracting each time loss, as illustrated in Figure 2.1 and definitions [2].
$\texttt{OT}-\texttt{SBT}$ $=\textrm{Loading Time}\ (\texttt{LT})$
$\texttt{LT}-\texttt{DT}$ $=\textrm{Operating Time}\ (\texttt{OpT})$
$\texttt{OpT}-\texttt{PLT}$ $=\textrm{Net Operating Time}\ (\texttt{NOpT})$
$\texttt{NOpT}-\texttt{QLT}$ $=\textrm{Valuable Time}\ (\texttt{VT})$ [2.1]
Moreover, the ratio between the production times can be used to define some well-known effectiveness indicators, which are enumerated in formulae [2]:
$\dfrac{\texttt{LT}}{\texttt{OT}}$ $=\textrm{Loading Rate}\ (\texttt{lo})$
$\dfrac{\texttt{OpT}}{\texttt{LT}}$ $=\textrm{Availability Rate}\ (\texttt{av})$
$\dfrac{\texttt{NOpT}}{\texttt{OpT}}$ $=\textrm{Performance Rate}\ (\texttt{pf})$
$\dfrac{\texttt{VT}}{\texttt{NOpT}}$ $=\textrm{Quality Rate}\ (\texttt{qu})$ [2.2]
@@
$\texttt{oee}=\texttt{av}\times\texttt{pf}\times\texttt{qu}$ [2.3]
@@
date date yyyy-$m$-$d$
start timestamp $h$:$m$:$s$
@@ 3.2 The discrete process
[…]
@@ 3.3 The continuous process
$\mathbf{y}_{n}|\mathbf{w}_{n}\sim N_{m}\left(\mathbf{u}_{n}\mathbf{H}_{u},\bm{\Sigma}_{u}\right)$ [3.2a]
$\mathbf{y}_{n}|c_{n}\sim N_{m}(\mathbf{v}_{n}\mathbf{H}_{v},\bm{\Sigma}_{v}),$ [3.2b]
where $m$ is the number of response variables; $\mathbf{H}_{u},\mathbf{H}_{v}$ denote coefficient matrices; $\bm{\Sigma}_{u},\bm{\Sigma}_{v}$ are covariance matrices; and $\mathbf{u}_{n}=\left[1\ \mathbf{w}_{n}\right]$, $\mathbf{v}_{n}=v(c_{n})$ are the covariate vectors at the $n$-th step, with $v(\cdot)$ a function of the hidden state. In particular, we propose considering the conditional expectation $v(c_{n})=\mathbb{E}\left[c_{n}|c_{n-1}\right]$, which, with the Dirichlet distribution assumptions, simplifies to the Dirichlet parameters in [3.1] normalized by their concentration parameters, i.e., $\mathbf{v}_{n}=\mathbf{a}/a_{0}$ or $\mathbf{v}_{n}=\bm{\alpha}_{c_{n-1}}/\alpha_{c_{n-1},0}$.
@@ 3.4 The adaptive algorithm
We explain the adaptive algorithm for the case of the parameters in distribution [3.2a]; the same procedure is applied to the parameters in [3.2b]. Let $\lambda\in(0,1]$ be a forgetting factor that accounts for the weight of past observations. As soon as a new sample $\mathbf{y}_{n}$ becomes available, the prior estimators $\mathbf{H}_{n-1},\bm{\Sigma}_{n-1}$ are updated to $\mathbf{H}_{n},\bm{\Sigma}_{n}$ (we omit the model subscripts in the parameters for clarity) through an adaptive algorithm described by the following equations:
$\gamma_{n}$ $=1+\lambda\gamma_{n-1}$ [3.3a]
$\mathbf{H}_{n}$ $=\mathbf{H}_{n-1}+\dfrac{\mathbf{P}_{n-1}\mathbf{u}_{n}^{T}}{\lambda+\mathbf{u}_{n}\mathbf{P}_{n-1}\mathbf{u}_{n}^{T}}\left(\mathbf{y}_{n}-\mathbf{u}_{n}\mathbf{H}_{n-1}\right)$ [3.3b]
$\bm{\Sigma}_{n}$ $=\bm{\Sigma}_{n-1}-\dfrac{1}{\gamma_{n}}\left[\bm{\Sigma}_{n-1}-\dfrac{\lambda\left(\mathbf{y}_{n}-\mathbf{u}_{n}\mathbf{H}_{n-1}\right)^{T}\left(\mathbf{y}_{n}-\mathbf{u}_{n}\mathbf{H}_{n-1}\right)}{\lambda+\mathbf{u}_{n}\mathbf{P}_{n-1}\mathbf{u}_{n}^{T}}\right]$ [3.3c]
$\mathbf{P}_{n}$ $=\dfrac{1}{\lambda}\left(\mathbf{P}_{n-1}-\dfrac{\mathbf{P}_{n-1}\mathbf{u}_{n}^{T}\mathbf{u}_{n}\mathbf{P}_{n-1}}{\lambda+\mathbf{u}_{n}\mathbf{P}_{n-1}\mathbf{u}_{n}^{T}}\right)$ [3.3d]
initialized with $\mathbf{H}_{0}=\bm{0},\ \bm{\Sigma}_{0}=\bm{0},\ \mathbf{P}_{0}=\mathbf{I}$ and $\gamma_{0}=0$$\gamma_{n}$ is the total weight of the sample. Note that if $\lambda=1$ then all the observations have the same weight and $\gamma_{n}=n$ is the sample size; if $\lambda<1$, past data gradually loses influence over time while recent data has a greater impact. Equations [3.3a]-[3.3d] can be obtained by extending to the multivariate case the Maximum-Likelihood-based proof provided in Alvarez et al.Alvarez et al. [2021], Theorem. 1. We refer to Appendix A for an alternative proof using a Bayesian approach.
@@ 3.5 Forecasting
After the training step, each distribution [3.2a]-[3.2b] produces a forecast of the responses. Later, these forecasts are combined using a minimum-variance criterion to obtain the final prediction [Roccazzella et al., 2022]. Once a new observation is available the update-prediction loop begins again. In particular, when the parameters are updated after the $n$-th observation is received we can write
@@
$\mathbb{E}_{n}\left[\mathbf{y}_{n+1}\right]\stackrel{{\scriptstyle def}}{{=}}\hat{\mathbf{y}}_{n+1}$ $=\mathbf{u}_{n+1}\mathbf{H}_{u}\mathbf{D}+\mathbf{v}_{n+1}\mathbf{H}_{v}(\mathbf{I}-\mathbf{D})$ [3.4a]
$\mathbb{V}\mathrm{ar}_{n}\left[\mathbf{y}_{n+1}\right]\stackrel{{\scriptstyle def}}{{=}}\hat{\bm{\Sigma}}_{n+1}$ $=\mathbf{D}\bm{\Sigma}_{u}\mathbf{D}+(\mathbf{I}-\mathbf{D})\bm{\Sigma}_{v}(\mathbf{I}-\mathbf{D}),$ [3.4b]
@@
$\delta_{j}=\dfrac{\sigma_{v,j}^{2}}{\sigma_{u,j}^{2}+\sigma_{v,j}^{2}},$ [3.5]
@@ 4.3 Clustering
Once the classification is complete, each sequence in the training set is segmented into labelled intervals corresponding to the detected operational states. These states are later used to learn the parameters of the discrete part of the model, as described in Section 3.2, starting from non-informative Jeffreys’ priors for the Dirichlet distributions [3.1], that is, $a_{k}=\alpha_{jk}=1/2,\ \forall j,k$.
@@ 4.4 Pseudo-code
$\bm{\Pi}$ $=\left\{\mathbf{a_{s}}=[a^{(\mathbf{s})}_{1}\ \ldots\ a^{(\mathbf{s})}_{K}],\mathbf{A_{s}}=\left(\alpha_{kj}^{(\mathbf{s})}\right)_{k,j\in\mathcal{K}},\ \mathbf{s}\in\mathcal{S}\right\}$
$\bm{\Psi}$ $=\left\{\mathbf{H}_{u}^{(\mathbf{s})},\bm{\Sigma}_{u}^{(\mathbf{s})},\mathbf{H}_{v}^{(\mathbf{s})},\bm{\Sigma}_{v}^{(\mathbf{s})},\ \mathbf{s}\in\mathcal{S}\right\}$
$\bm{\Omega}$ $=\left\{\mathbf{P}_{u}^{(\mathbf{s})},\gamma_{u}^{(\mathbf{s})},\mathbf{P}_{v}^{(\mathbf{s})},\gamma_{v}^{(\mathbf{s})},\ \mathbf{s}\in\mathcal{S}\right\}$
@@
2:  $\bm{\Pi}$ $\triangleright$ Dirichlet parameters
3:  $\bm{\Psi}$ $\triangleright$ model parameters
4:  $\bm{\Omega}$ $\triangleright$ state parameters
5:  $\lambda_{u},\lambda_{v}$ $\triangleright$ forgetting factors
6:  $\mathbf{x}_{n}=\left[\mathbf{z}_{n}\ \mathbf{w}_{n}\right]$ $\triangleright$ covariates
7:  $c_{n-1},c_{n},\mathbf{y}_{n}$ $\triangleright$ state labels, responses
@@
16:  Update $\gamma_{u}^{(\mathbf{s})},\mathbf{H}_{u}^{(\mathbf{s})},\bm{\Sigma}_{u}^{(\mathbf{s})},\mathbf{P}_{u}^{(\mathbf{s})}$ using equations [3.3a], [3.3b], [3.3c], [3.3d] respectively
17:  Update $\gamma_{v}^{(\mathbf{s})},\mathbf{H}_{v}^{(\mathbf{s})},\bm{\Sigma}_{v}^{(\mathbf{s})},\mathbf{P}_{v}^{(\mathbf{s})}$ using equations [3.3a], [3.3b], [3.3c], [3.3d] respectively
@@
2:  $\bm{\Psi}$ $\triangleright$ model parameters
3:  $\left\{\mathbf{o}_{1},\ldots,\mathbf{o}_{K}\right\}$ $\triangleright$ centroids
4:  $\mathbf{t}_{n}$ $\triangleright$ classification variables
5:  $\mathbf{x}_{n+1}=\left[\mathbf{z}_{n+1}\ \mathbf{w}_{n+1}\right]$ $\triangleright$ covariates
@@
[…]
@@
[…]
14:Compute $\mathbf{D}$ using equation [3.5]
15:Compute $\hat{\mathbf{y}}_{n+1}$ using equation [3.4a]
16:Compute $\hat{\bm{\Sigma}}_{n+1}$ using equation [3.4b]
@@ 5 Application to a real case study
The time variables OpT and NOpT are selected as response variables due to their strong correlation (0.83), which fully justifies the use of a multivariate model. According to the scheme presented in Figure 2.1 and the definitions [2], [2] and [2.3], predicting these variables allows for the computation of time losses due to low production speed and the performance index. Additionally, in scenarios where time measurements are equispaced and scheduled stops are predetermined (i.e., with deterministic OT, SBT and LT), these predictions also enable the estimation of time losses due to unexpected stops and the availability index.
@@ Appendix A Appendix
[…]
@@
    $N\tilde{S}$ $=\left(Y-X\tilde{B}\right)^{T}\left(Y-X\tilde{B}\right)+\left(\tilde{B}-B_{0}\right)^{T}V_{0}\left(\tilde{B}-B_{0}\right).$ [A.1a]
@@
    $\mathbb{E}\left[B|\Lambda\right]$ $=\tilde{B}=B_{0}+\left(X^{T}X+V_{0}\right)^{-1}X^{T}\left(Y-XB_{0}\right)$ [A.2a]
    $\mathbb{E}\left[\Lambda|Y\right]$ $=(N_{0}\underset{\underset{\text{prior mean}}{\uparrow}}{\Lambda_{0}}+N\tilde{S})/(N_{0}+N)$
    $=\Lambda_{0}-\dfrac{N\left(\Lambda_{0}-\tilde{S}\right)}{N_{0}+N}.$ [A.2b]
@@
    $Y\$$=\ \mathbf{y}_{n}\ \text{ (new responses, row vector $1\times m$)}$
    $X\$$=\ \mathbf{u}_{n}\ \text{ (new predictors, row vector $1\times p$)}$
    $\tilde{B}\$$=\ \mathbf{H}_{n}\ \text{ (posterior mean, $p\times m$)}$
    $B_{0}\$$=\ \mathbf{H}_{n-1}\ \text{ (prior mean, $p\times m$)}$
    $V_{0}\$$=\ \lambda\mathbf{P}_{n-1}^{-1}\ \text{ (prior state matrix, $p\times p$)}$
    $(X^{T}X+V_{0})^{-1}\$$=\mathbf{P}_{n}\ \text{ (posterior state matrix, $p\times p$)}$
    $\gamma_{n}\$$=1+\lambda+\cdots\lambda^{n-1}\ \text{(scalar)}$
@@
    we obtain equation [3.3d]. For the posterior mean $\tilde{B}=\mathbf{H}_{n}$, using expression [A.2a] we have
@@
    yielding equation [3.3b]. On the other hand, from the definition [A.1a] we get
@@
    $N_{0}\$$=\ \lambda\gamma_{n-1}$
    $N\$= $=\ 1$
    $\Lambda_{0}\$$=\bm{\Sigma}_{n-1}\ \text{ (prior mean variance, $m\times m$)}$
    $\mathbb{E}\left[\Lambda|Y\right]\$$=\ \bm{\Sigma}_{n}\quad\text{ (posterior mean variance, $m\times m$)}$
    and using expression [A.2b]
    $\bm{\Sigma}_{n}=\mathbb{E}\left[\Lambda|Y\right]\$$=\Lambda_{0}-\dfrac{N\left(\Lambda_{0}-\tilde{S}\right)}{N_{0}+N}$